5 divided by x minus 6 divided by y equal 1
17 divided by x plus 30 divided by y equals 16

Answer by anshu_a(16)

5 divided by x minus 6 divided by y equal 1
17 divided by x plus 30 divided by y equals 16
sol :
5/x - 6/y = 1-------(1)
17/x + 30/y =16 -----(2)
multyiply (1) by 5 we will get
25/x - 30/y = 5 ----------(3)
17/x + 30/y = 16 ---------(2)
add (2) and (3) we will get
25/x+17/x = 21
42/x = 21 or x = 2
put value of x in (1) we will get
5/2 -6/y = 1 or 6/y = 5/2 - 1
6/y = 3/2 or y = 2/3*6
y = 4